OpenStudy (anonymous):

Best response if you can help: Line p passes through the point (5, 1) and has a slope of -4/5. What is the equation of line p in standard form? Hint: Point slope form of of an equation of a line y-y1= m(x-x1) a. 4x + 5y = 25 b. 5x + 4y = 25 c. 5x + 4y = 29 d. 4x + 5y = 29

OpenStudy (anonymous):

its B

OpenStudy (anonymous):

Point slope form \[(y -y_1) = m* (x-x_1)\] y1 and x1 are points and m is the slope

OpenStudy (anonymous):

I got it anyways ty ;p

OpenStudy (anonymous):

yipp I realized haha ty tho

OpenStudy (anonymous):

I actually got 4x + 5y = 25 tho...